OpenStudy (anonymous):

Can someone quickly help me with 4 questions!

OpenStudy (anonymous):

Are they easy ?

OpenStudy (anonymous):

It depends and two of them are going to be written responses but since I don't have a laptop right now i'm using my older sisters but she has to take a college test and be done by twelve so I have to hurry

OpenStudy (anonymous):

You want me to show you the questions

OpenStudy (anonymous):

of course lol

OpenStudy (anonymous):

That'll be nice

OpenStudy (anonymous):

Loll sorry okay let me post them here

OpenStudy (anonymous):

The graph below is represented by which system of inequalities?

OpenStudy (anonymous):

x + y > 4 and 2x − y < 10 x + y < 4 and 2x − y < 10 x + y > 4 and 2x − y > 10 x + y < 4 and 2x − y > 10

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

Oh yeah, I am not good with this, I will wait till' the next one :))

OpenStudy (anonymous):

That's fine :)

OpenStudy (anonymous):

x + y > 4 and 2x − y < 10

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Next question

OpenStudy (anonymous):

Welcum and thanks for the medal

OpenStudy (anonymous):

Np! sorry my computer is going slow

OpenStudy (anonymous):

NP!

OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 124 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 28, how many women auditioned? 40 44 70 80

OpenStudy (anonymous):

We will use system of equations here

OpenStudy (anonymous):

M = number of men W = number of women Use a system of equations: M + W = 124 2M - 3W = 28 change M + W = 124 to: M = 124 - W plug that into 2M - 3W = 28 and solve for W 2(124 - W) - 3W = 28 248 - 2W - 3W = 28 -5W = -220 W = 44 44 women auditioned

OpenStudy (anonymous):

2nd answer

OpenStudy (anonymous):

Thank you ok next are 1 out of the 2 written answers

OpenStudy (anonymous):

For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (anonymous):

Part 1: You are buying potatoes and meat. If you buy 2 pounds of potatoes and one pound of meat it will cost you $10, if you buy 3 lbs of potatoes and 4 lbs of meat you will pay $25. How much do potatoes and meat cost? Part 2: Let x be the cost of potatoes and y the cost of meat (in dollars per pound). Then 2 lbs of potatoes and 1 lb of meat will cost 2x+y=10 and 3 lbs of potatoes and 4 lbs of meat will cost 3x+4y=25. Solving the equations: from the first equation we get y=10-2x. Plugging in the second gives 3x+4(10-2x) = 25. Getting rid of parentheses we get -5x + 40 = 25 and therefore 5x=15, giving us x=3. Therefore y=10-2*3=4. Potatoes cost $3 per pound and meat $4.

OpenStudy (anonymous):

Ok perfect @UditKulka ready for the next one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

] You graph a system of equations and the two lines overlap, or lie right on top of one another. Explain what this tells you about the: solution to the system the two equations the result if the system was solved algebraically

OpenStudy (anonymous):

if they intercept at one point, the solution is the intersecting point if they are parallel , and do not intercept, there is no solution if they are parallel and they are same line, there is infinite solution where the lines intersect, the equations are equal. (If they were simultaneous equations, these would be the solutions) where the lines lie on top of each other, one equation can be rearranged to be the other equation precisely. this means when solving, you would get y=y and x=x. there are infinite solutions in this case (every point along the line would be a solution). (as an addition: where the lines dont cross at all, there is no real solution to the system (e.g. x^2 = x^2 + 5))

OpenStudy (anonymous):

@UditKulka I can't use the first one because my teacher would say its against the academic integrity rules to copy Can you please just make one for me

OpenStudy (anonymous):

It would be my pleasure to help u but i have to sit for a english test sry

OpenStudy (anonymous):

Oh okay thanks for the other questions though!

OpenStudy (anonymous):

No problem